Operations Report – May 10th

Crew 212 Operations Report 10-MAY-2019

SOL: 5

Name of person filing report: Paolo Guardabasso

Non-nominal systems: None

Notes on non-nominal systems: Nothing to report

Generator (hours run last night): [From Director]

On (last night): 9:00 PM

Off (this morning): [From Director]

Solar— SOC 99% at 7:40 PM

Diesel Reading – % [From Director]

Propane Reading – % [From Director]

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 200 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 155 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143454.5

Toilet tank emptied: No

Deimos rover used: no (out of order)

Hours:
Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes

Hours: 92.6h

Beginning charge: 100%

Ending charge: 57%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 59.1h

Beginning charge: 100%

Ending charge: 74%

Currently charging: No

Curiosity rover used: Yes

Hours: 94.7h

Beginning charge: 100%

Ending charge: 75%

Currently charging: Yes

Notes on rovers: Spirit showed an anomalous power consumption, hitting 70% charge after about 10 minutes of driving. It forced the EVA crew to come back to the Hab, where the charge was 57%
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report

CrewCar used and why, where? Off Campus

General notes and comments: We received the visit from two journalists, who toured the station and interviewed the crew members. One EVA was performed this morning and the crew was accompanied by two journalists. The EVA in the afternoon postponed because of bad weather, but it was successfully carried out 30 minutes later.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: It was depressurized to allow the EVA crew to take the drones for the experiments.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: a crew member had a minor injury.

Questions, concerns, and requests to Mission Support: Following the issue of the heater, we were aware of the procedure in the handbook and we already tried keeping the hot water sink fully on for more than 10 seconds, but the heater still didn’t activate. Both gas and water are at 100%, and the propane and water lines valves are open.